OpenStudy (anonymous):

Please help me D: Write an equation of the line that is perpendicular to 2x + 4y = 6 and passes through the point (6,4).

OpenStudy (anonymous):

what's the slope of the given line, 2x + 4y = 6 ????

OpenStudy (anonymous):

the slope of that like would be -1/2 then you find the opposite reciprocal of that and it would be 2 so then you take the point-slope formula which is y-y1=m(x-x1) and plug in those numbers so it would be y-4=2(x-6) then you solve for y!

OpenStudy (anonymous):

or u could just leave it in point slope form (as is).......

OpenStudy (anonymous):

Thank you both very much c: The help was greatly appreciated. ^-^

OpenStudy (anonymous):

welcome!